# Whose Boat Does it Float? Improving Personalization in Preference Tuning via Inferred User Personas

### v1

@@ 5.3 DPO Sans Personas is Not All You Need
5.4 PT Personalizes to User-Specified Needs
To verify PT aids user-written needs, we run a pilot study with 10 CS students who often use LLMs. We pick 12 queries from BeaverTails and Anthropic HHH (party plans, job search) that students may ask LLMs Zhao et al. (2024b). Users first write three personas per prompt so models must adapt to personas, and then rate $\textsc{PT}_{\textsc{dpo}}$ and DPO outputs—using prompts and their personas as inputs—on 1-5 on Answerability (answering queries) and Personalization (tailoring to personas). This mirrors our offline metrics (section 4.3) and assesses whether models meet the needs of users who specified them.
Both models have high answerability but $\textsc{PT}_{\textsc{dpo}}$ is significantly more personalized (Figure 6), so $\textsc{PT}_{\textsc{dpo}}$ better supports user-written needs. As users perceive $\textsc{PT}_{\textsc{dpo}}$ as more helpful, future work can assess how PT impacts long-term trust or engagement Serino et al. (2005) and if it aids downstream tasks Balepur et al. (2024b); Mozannar et al. (2024).
5.5 Promises and Pitfalls of Persona Tailoring
We show strengths and issues of $\textsc{PT}_{\textsc{dpo}}$ (Figure 7). $\textsc{PT}_{\textsc{dpo}}$ alters answers for information needs and tailors mnemonics to learning styles (blue), showing its promise for downstream personalization tasks. However, $\textsc{PT}_{\textsc{dpo}}$ assumes personas are harmless, leading to sycophancy Sharma et al. (2024); adversaries can exploit this and use personas to get inaccurate, biased, or irrelevant text (red). To solve this, future work can create undesired personas to teach LLMs when to abstain Wang et al. (2024b).
@@ 6 Related Work
Personalization:
LLM personalization aims to produce text aligned with a user’s needs, interests, or attributes Zhang et al. (2024c); Kirk et al. (2024a); Chen et al. (2024b). Personalization can aid engagement Pardini et al. (2022), trust Serino et al. (2005), and learning Bernacki et al. (2021) across recommender systems Islam et al. (2014), gaming Tondello et al. (2017), chatbots Madotto et al. (2019), and tutoring frameworks Park et al. (2024).
@@
For personalization, many works use personas—textual user descriptions Salminen et al. (2020)—in prompts Jandaghi et al. (2024); Liu et al. (2024a); Deshpande et al. (2023), but we train LLMs with personas. Stephan et al. (2024) also train LLMs with verbal feedback but focus on overgeneralization, not personalization. Lee et al. (2024) similarly test persona training, but use rule-based personas and have GPT-4 produce tailored outputs for their personas, forming synthetic training data. In contrast, we infer personas in preference data, which has more flexibility and does not require a teacher model that already excels in personalization; the latter is key in specific domains like education where we cannot rely on LLMs as ground-truth, but user preference data can be curated Liu et al. (2024c).
@@
Lastly, works have tried eliciting personas from a user’s interaction history Li et al. (2023); Handa et al. (2024); Jin et al. (2024). Our PI task is a form of preference elicitation, deriving personas from a single pairwise comparison (section 2). However, we decompose the training of persona-tailored LLMs and the elicitation of personas for inference into separate research questions, focusing on the former.
Preference Subjectivity:
@@
User preferences are subjective and depend on a user’s opinions, traits, and values Bakker et al. (2022); Kirk et al. (2024b); Agnew et al. (2024). As a result, researchers have studied social choice theory Conitzer et al. (2024) and pluralistic alignment Sorensen et al. (2024); Liu et al. (2024a) to handle subjectivity, executed via Bayesian preference tuning Yang et al. (2024); Handa et al. (2024), merging models individually aligned to groups Chakraborty et al. (2024); Hwang et al. (2023), multi-LLM collaboration Feng et al. (2024); Wu et al. (2024b) and multi-objective reward modeling Wang et al. (2024a); Zeng et al. (2024).
@@
While these methods capture preferences across examples, we infer preferences in single pairwise comparisons via abductive reasoning—finding evidence to explain outcomes Peirce (1974). This reasoning has been used in commonsense Zhao et al. (2023), question answering Wu et al. (2024a), and reading comprehension Du et al. (2021) tasks, but we use it to infer personas of users that prefer chosen or rejected outputs. Very recent and concurrent work also explores how contexts can change preferences on responses to improve human evaluation Malaviya et al. (2024) and preference modeling Pitis et al. (2024), but we are the first to treat these contexts as personas to tailor responses for users.
@@ 7 Conclusion
Abductive reasoning—explaining when outputs are preferred—greatly aids personalization via persona inference (PI) and tailoring (PT). PI infers uncommon but valid reasons to prefer rejected responses which form harder evaluations, so future work can curate better, long-tail user needs Yin et al. (2012) to stress-test personalization. Users favor PT’s outputs, but there are many questions to study, like if PT helps users finish tasks Mozannar et al. (2024) and if users can easily verbalize needs as prompts Taylor (1962). Despite PT’s success, we may need debiasing and abstention Meade et al. (2022); Wang et al. (2024b) to curb harmful or biased personas.
Beyond personalization, personas reveal implicit differences in chosen and rejected outputs, which can be further surfaced via contrastive topic models Zhong et al. (2023). Our personas convey user traits and needs, but other types—like moral arguments Rao et al. (2023), cultural values Kirk et al. (2024b), and authorship Wang et al. (2023)—could adapt LLMs to diverse constitutions, cultures, and writing styles. Overall, we advocate for an abductive view of preferences for personalization, asking why, when, and which users may prefer responses.
@@ 8 Limitations
One limitation is that our persona tailoring method assumes there is always an input persona that is relevant to the prompt. However, as we show in section 5.5, a user could give personas that are irrelevant to the prompt, harming our model outputs. To address this, future work could explore producing intentionally irrelevant personas and training the model to abstain on them. Further, to address cases where a specific persona is not needed as input, researchers could route prompts to a model trained without personas, or use a default persona like a general system prompt; we test the latter in Appendix A.9.
We also acknowledge that we cannot fully evaluate all possible personas and user needs that a model must cater to. To enhance persona diversity for evaluation, we use personas derived from both chosen and rejected responses in offline evaluation (section 5.3), resulting in two diverse user needs per input prompt. Further, in our pilot study, we ask annotators to write three different personas per input prompt, and we find few exact matches in personas. We believe there is promising future work to rigorously test how personalization impacts diverse users, including if personalization harms or helps users across cultures Kirk et al. (2024b), if personalization truly helps users in downstream tasks Mozannar et al. (2024), and if users can accurately articulate their preferences Handa et al. (2024).
Finally, LLMs are sensitive to prompt variations Sclar et al. (2024), so our persona inference and tailoring results may differ by prompt design. To address this, we follow best practices in prompt engineering Schulhoff et al. (2024) to ensure each model has a fair opportunity to generate valid personas in persona inference. Further, due to resource constraints, we only experimented with a single combination of student and teacher models for persona tailoring. We believe future work can investigate the potential of using the same model for both persona generation and training, employing user personas in a self-evolving learning framework.
@@ 9 Ethical Considerations
Personalization can raise ethical concerns when using personas tied to sensitive attributes like race or gender, which risks perpetuating stereotypes Kantharuban et al. (2024). As a result, we did not focus on these types of personas in this work, and instead studied high-level interests, personality traits, and information needs (section 2). While users specifying sensitive personas directly is less concerning, issues arise when practitioners use preference elicitation techniques Li et al. (2023) to infer such personas from past interactions and employ them for personalization. Although we observed no instances of protected attributes in persona inference, our model trained on “safe” personas could still produce harmful responses if prompted to (section 5.5). We recommend future works to explore safeguards, such as classifiers to flag harmful personas pre-inference, and conduct user studies to understand which types of personas users prefer seeing reflected in outputs, guiding responsible personalization efforts.
Further, personalization could risk selection or confirmation bias Hernán et al. (2004); Klayman (1995). For example, if a user already has a particular view and requests information that is aligned with that view, our model will provide a response that confirms this user’s view, which may be harmful in cases such as misinformation Zhou and Shen (2022). There is a tradeoff between generating personalized and balanced responses, and researchers can explore future task setups that encourage generating balanced outputs Zhang et al. (2024b) to better quantify the effects of this tradeoff.
Acknowledgments
We would like to thank the CLIP lab at the University of Maryland and our external collaborators for their feedback. We also thank Dang Nguyen, Yu Hou, Connor Baumler, Paiheng Xu, Zongxia Li, Feng Gu, Naina Balepur, and Atrey Desai for their annotation help. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber), iis-2339746 (Rudinger), and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research.
@@ A.4 LLM Judge Details
To assess GPT-4o’s reliability, we calculate the agreement of one Ph.D. student (an author) with the LLM in persona inference accuracy (section 3.1). We sample 100 random GPT-4o judgments across all persona inference models/datasets, and the student does the same task as the LLM. We find a high raw agreement of 86%, so the metric is reliable.
@@ A.7 Extended Ablations
We present further results and descriptions of our ablation studies (section 5.2). We first explore length discrepancies in model outputs (Table 7), finding that models using different generation strategies (center) have large length discrepancies over one sentence long, while models using the same generation strategy (top) have similar lengths. To avoid verbosity bias in ablations Zheng et al. (2024), we only use pairs with the same number of sentences. The comparison between DPO and $\textsc{PT}_{\textsc{dpo}}$ in section 5.3 also has discrepancies in sentence count, but as this is a personalization comparison and personas often relate to length (e.g. “The user is comprehensive”), we feel length adjustments are not appropriate.
Another interesting finding in our ablations is that on BeaverTails, supervised fine-tuning with personas does not surpass few-shot prompting with personas. To give a potential explanation, we also conduct ablations of models trained without personas (Table 3). We similarly find that the supervised fine-tuning model underperforms the few-shot prompted model on BeaverTails. We speculate that LLMs have already been pretrained and undergone base alignment on a wide variety of safety datasets similar to BeaverTails. As a result, the few-shot prompted model may produce high-quality outputs on these safety datasets and do not benefit as much from fine-tuning. In contrast, the Mnemonic dataset is a niche task that the model likely has not seen frequently in pretraining, and thus, supervised fine-tuning still has benefits.
@@ A.11 Repetitive Text on Anthropic HHH
On Anthropic HHH, our manual analysis found that models trained on this dataset and using greedy decoding could repeat text without generating the end-of-text token (around $20\%$ of cases). For example, when given a user request for celery recipes, the DPO model produces the text: “You can eat it raw, add it to salads, or use it in soups and stews. You can also make celery juice, or use it in a celery juice cocktail. You can also use it in a celery juice smoothie, or in a celery juice cocktail…” until the maximum token length is reached; some outputs in the dataset have this repetitive nature, which we believe could lead to this behavior. To avoid re-running all experiments by filtering these examples in training data, we omitted these outputs in inference to ensure no model had a fair/unfair advantage in generating these types of repetitive outputs.
To ensure these outputs do not impact our findings, we run our evaluations on Anthropic HHH using the full test set. We repeat our experiments in section 5.1 with this constraint in Table 15, and our experiments in section 5.3 in Table 16. Our findings are consistent (and in fact, stronger), showing that PT typically has net improvements in personalization while maintaining quality, while PT excels in personalizing to the uncommon but valid needs linked to rejected responses. Thus, our results confirm that PT is still a strong technique despite this issue.
In future directions of this work, we plan to explore varied decoding strategies (e.g. greedy decoding, beam search, nucleus sampling, repetition penalties) to investigate how they impact model personalization Zarrieß et al. (2021).
@@ A.12 Annotator Instructions
We present instructions given to annotators when writing personas and assessing outputs for prompts and personas in Figure 9 and Figure 10, respectively. The annotators are four graduate students and one undergraduate student studying computer science and have research experience with LLMs, natural language processing, and information retrieval; none of them are authors. We did not collect any personal information about the participants, except for their preferences in personas. Participants were rewarded with gift cards and purchased meals which averaged around \$20 per hour, above our region’s minimum wage. Our setup was approved by an Institutional Review Board to mitigate any potential harms. Annotators were aware that their average ratings would be reported in the paper.
@@
BeaverTails Anthropic HHH Mnemonic
Table 15: Win, tie, and loss rates of generation methods (FS, SFT, DPO) with and without personas $\mathcal{P}$ in pairwise comparisons from the Prometheus judge when using the full Anthropic HHH dataset. Our results are even stronger than in Table 2, with even larger $\Delta\textsc{PQ}$ increases.
@@
Table 16: Comparison of personalization abilities of DPO and $\textsc{PT}_{\textsc{dpo}}$ when using the full Anthropic HHH dataset. Our results are even stronger than in Table 4; DPO always bests $\textsc{PT}_{\textsc{dpo}}$ in personalization and response quality.

### v2

@@ 6 Related Work
Below, we review research on LLM personalization (section 6.1) and preference subjectivity (section 6.2)—showing how they relate to our strategy of inferring personas (section 2) and training models to tailor to them (section 4).
6.1 Personalization
LLM personalization aims to steer models toward user-specific requests, interests, and values Zhang et al. (2024c); Chen et al. (2024b), improving user trust Serino et al. (2005), engagement Pardini et al. (2022), and learning Bernacki et al. (2021).
@@ 6.1 Personalization
For personalization, many works use personas—textual user descriptions Salminen et al. (2020)—in prompts Jandaghi et al. (2024); Liu et al. (2024a); Deshpande et al. (2023), but we train LLMs with personas. Stephan et al. (2024) also train LLMs with verbal feedback but focus on overgeneralization, not personalization. Lee et al. (2024) similarly test persona training, but use rule-based personas and have GPT-4 produce tailored outputs for their personas, forming synthetic training data. In contrast, we infer personas in preference data, which has more flexibility and does not require a teacher model that already excels in personalization; the latter is key in specific domains like education where we cannot rely on LLMs as ground-truth, but user preference data can be curated Liu et al. (2024c). We show this empirically via large personalization gains on our mnemonic dataset (Table 2, Table 4). Further, unlike these works, we run a comprehensive evaluation, covering persona accuracy (section 3.1), plausibility, and harmfulness (section 3.3), as well as response quality with a small user study (section 5.4).
@@
Lastly, works have tried eliciting personas from a user’s interaction history Li et al. (2025); Handa et al. (2024); Jin et al. (2024). Our PI task is a form of preference elicitation, deriving personas based on pairwise comparisons (section 2). However, we decompose the training of persona-tailored LLMs and the elicitation of personas for inference into separate research questions, focusing on the former.
6.2 Preference Subjectivity
@@
User preferences are subjective and depend on a user’s opinions, traits, and values Bakker et al. (2022); Kirk et al. (2024); Agnew et al. (2024). As a result, researchers have studied social choice theory Conitzer et al. (2024) and pluralistic alignment Sorensen et al. (2024); Liu et al. (2024a) to capture subjectivity, executed with Bayesian preference training Yang et al. (2024); Handa et al. (2024), combining models individually aligned to groups Chakraborty et al. (2024); Hwang et al. (2023), multi-LLM collaboration Feng et al. (2024); Wu et al. (2025), and multi-objective reward modeling Wang et al. (2024a); Zeng et al. (2024).
@@
Complementing these approaches, we design PI, a method that can infer preferences from pairwise comparisons via abductive reasoning—finding evidence to explain outcomes Peirce (1974). This reasoning has been used in commonsense Zhao et al. (2023), question answering Balepur et al. (2024b), and reading comprehension Du et al. (2021) tasks, but we use it to infer personas of users that prefer chosen or rejected outputs. Very recent and concurrent work also explores how contexts can change preferences on responses to improve human evaluation Malaviya et al. (2024) and preference modeling Pitis et al. (2024), but we are the first to treat contexts as personas to tailor responses for users.
@@ 7 Conclusion
Abductive reasoning—explaining when outputs are preferred—greatly aids personalization via persona inference (PI) and tailoring (PT). PI infers uncommon but valid reasons to prefer rejected responses which form harder evaluations, so future work can collect more long-tail user needs Yin et al. (2012) to stress-test personalization. Notably, our method generalizes to real user needs without curating potentially sensitive data, showing LLM-inferred personas can improve personalization with less privacy concerns Peng et al. (2024). Users prefer PT’s outputs, but there are remaining questions: can users faithfully verbalize their needs for prompts Taylor (1962) and can PT help users finish tasks Mozannar et al. (2025)? Despite PT’s success, we may need debiasing and abstention Meade et al. (2022); Wang et al. (2024c) to curb harmful or biased personas.
Beyond personalization, personas reveal implicit differences in chosen and rejected outputs; we surface these via persona token saliency, but tools like contrastive topic models may provide finer-grained differences Zhong et al. (2023). Our personas convey user needs, but other types—like moral arguments Rao et al. (2023), cultural values Kirk et al. (2024), authorship Wang et al. (2023), and knowledge levels Shu et al. (2024)—could adapt LLMs to varied constitutions, cultures, writing styles, and user expertise. Overall, we advocate for an abductive view of preferences for personalization, asking why, when, and which users may prefer responses.
Acknowledgments
We thank the clip lab at the University of Maryland and our external collaborators for their feedback. Specifically, we wish to thank Dang Nguyen, Yu Hou, Connor Baumler, Paiheng Xu, Zongxia Li, Feng Gu, Naina Balepur, and Atrey Desai for annotation help. We are also grateful to Zichao Wang, Vivek Iyer, Brihi Joshi, Robin Jia, John Lalor, and Nischal Kumar for insightful discussions. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber), iis-2339746 (Rudinger), and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research.
@@ 8 Limitations
One limitation is our PI strategy uses just one example to infer personas. In some cases, this may pose the risk of overfitting, but with our LLMs, we find personas rarely overfit to the prompt and response (section 3.3). Further, in some tasks, we may desire more than one example to find more nuanced personas. For instance, from multiple examples, we could infer users prefer short answers for queries on news and long answers for queries on food. Such a task can further challenge LLMs in abductive reasoning and could lead to personalization systems that consider personas from diverse angles. We find it extremely promising that our strategy of PI with just one example yields large personalization gains (Figure 6), and we hope future work can extend our PI task to capture diverse, multi-aspect personas.
Our persona tailoring method also assumes there is always an input persona that is relevant to the prompt. However, as we show in section 5.5, a user could provide input personas that are irrelevant to the prompt, degrading our model outputs. To address this, future work could explore producing intentionally irrelevant personas and training the model to abstain on them. Further, to address cases where a specific persona is not needed as input, researchers could route prompts to a model trained without personas, or use a default persona like a general system prompt; we test the latter in Appendix A.9.
We also note that we cannot capture all possible personas a model must cater to. For diversity, we use personas derived from both chosen and rejected responses in offline evaluation (section 5.3), resulting in two diverse user needs per input prompt. Further, in our pilot study, we ask annotators to write three unique personas per input prompt, and we find few exact matches. We believe there is promising future work in testing how personalization impacts diverse users, including how personalization helps users across cultures Kirk et al. (2024), if personalization truly helps users in downstream tasks Mozannar et al. (2025), and if users accurately articulate their specific preferences Handa et al. (2024).
Lastly, as is true in any machine learning model, there are tradeoffs between efficiency and performance; $\textsc{PT}_{\textsc{dpo}}$ produces the most personalized responses (section 5.2), but requires the most resources to train. Regardless, PT largely enhances personalization with every generation method (i.e., Prompting, SFT, and DPOsection 5.1), showing we can accommodate practitioners with varying resource budgets.
@@ 9 Ethical Considerations
Personalization can raise ethical concerns when using personas tied to sensitive attributes like race or gender, which risks perpetuating biases Hou et al. (2025) and stereotypes Kantharuban et al. (2024). Thus, we instead study high-level interests, personality traits, and needs (section 2). While users specifying sensitive personas directly is less concerning, issues arise when practitioners use preference elicitation techniques Li et al. (2025) to infer such personas from past interactions and apply them for personalization. Although we observed no instances of protected attributes in persona inference, our model trained on “safe” personas could still produce harmful responses if prompted to (section 5.5). We urge future works to explore safeguards like classifiers to flag harmful personas pre-inference, and conduct user studies to understand which types of personas users prefer seeing reflected in outputs, informing responsible personalization efforts.
Further, personalization could risk selection or confirmation bias Hernán et al. (2004); Klayman (1995). For example, if a user already has a particular view and requests information that is aligned with that view, our model will provide a response that confirms this user’s view, which may be harmful in cases such as misinformation Zhou and Shen (2022). There is a tradeoff between producing personalized and balanced responses, and researchers can explore future task setups that encourage generating balanced outputs Zhang et al. (2024b); Balepur et al. (2025b) to study its effects.
@@ A.4 LLM Judge Details
To assess GPT-4o’s reliability, we calculate the agreement of three Ph.D. students (two authors, one external) with the LLM in persona inference accuracy (section 3.1). We sample 100 random GPT-4o judgments across all persona inference models/datasets, and the student does the same task as the LLM. We find a high raw agreement of 90%, so the metric is reliable. The annotators have a Fleiss’ $\kappa$ inter-annotator agreement of $0.59$ Fleiss et al. (1981), indicating moderate to substantial agreement.
@@ A.7 Extended Ablations
We present further results and descriptions of our ablation studies (section 5.2). We first explore length discrepancies in model outputs (Table 7), finding that models using different generation strategies (center) have large length discrepancies over one sentence long, while models using the same generation strategy (top) have similar lengths. To avoid verbosity bias in ablations Zheng et al. (2024a), we only use pairs with the same number of sentences. The comparison between DPO and $\textsc{PT}_{\textsc{dpo}}$ in section 5.3 also has discrepancies in sentence count, but as this is a personalization comparison and personas often relate to length (e.g. “The user is comprehensive”), we feel length adjustments are not appropriate.
Another interesting finding in our ablations is that on BeaverTails, supervised fine-tuning with personas does not surpass few-shot prompting with personas. To give a potential explanation, we also conduct ablations of models trained without personas (Table 3). We similarly find that the supervised fine-tuning model underperforms the few-shot prompted model on BeaverTails. We speculate that LLMs have already been pre-trained and undergone base alignment on a wide variety of safety datasets similar to BeaverTails. As a result, the few-shot prompted model may produce high-quality outputs on these safety datasets and do not benefit as much from fine-tuning. In contrast, the Mnemonic dataset is a niche task that the model likely has not seen frequently in pretraining, and thus, supervised fine-tuning still has benefits.
@@ A.11 Repetitive Text on Anthropic HHH
On Anthropic HHH, our manual analysis found that models trained on this dataset and using greedy decoding could repeat text without generating the end-of-text token (around $20\%$ of cases). For example, when given a user request for celery recipes, the DPO model produces the text: “You can eat it raw, add it to salads, or use it in soups and stews. You can also make celery juice, or use it in a celery juice cocktail. You can also use it in a celery juice smoothie, or in a celery juice cocktail…” until the maximum token length is reached; some outputs in the dataset have this repetitive nature, which we believe could lead to this behavior.
To ensure these outputs do not impact our findings, we run our evaluations on Anthropic HHH using a subset of the test set where we filter these repetitive model outputs. We repeat our experiments in section 5.1 with this constraint in Table 16, and our experiments in section 5.3 in Table 15. Our findings are consistent, showing that PT typically has net improvements in personalization while maintaining quality, while PT excels in personalizing to the uncommon but valid needs linked to rejected responses. Thus, our results confirm that PT is still a strong technique despite this issue.
@@ A.12 Annotator Instructions
Apart from our LLM judge agreement evaluation, we have users assess the quality of personas (section 3.3) and the quality of personalized outputs (section 5.5). We discuss the protocols for both of these studies here.
The dimensions and instructions given to users for assessing persona quality are in section 3.3. For applicability, annotators had an ordinal Krippendorff’s $\alpha$ agreement of $0.40$, indicating moderate agreement and highlighting the subjective nature of applicability. For the other binary metrics of plausibility, offensiveness, and overfitting, all three users agreed on the gold label in $95\%$, $99\%$, and $94\%$ of cases.
We present instructions given to annotators when writing personas and assessing outputs for prompts and personas in Figure 9 and Figure 10, respectively. Our annotators have varied educational backgrounds, pursuing undergraduate, master, and doctoral degrees, and study diverse research fields including NLP, machine learning, HCI, information science, education, linguistics, and social networks; all of them have experience with using LLMs in either for their personal or professional needs.
For both of the qualitative analyses, we did not collect any personal information about the participants. Participants were rewarded with gift cards averaging \$20 per hour, above our region’s minimum wage. Our setup was approved by an Institutional Review Board to mitigate any potential harm. Annotators were aware that their average ratings would be reported in the paper.
@@
HHH Chosen DPO+$\mathcal{P}_{retr}$ PT+$\mathcal{P}_{retr}$ 37.2/22.6/40.2 32.6/21.3/46.0 -10.4
DPO+$\mathcal{P}_{gold}$ PT+$\mathcal{P}_{gold}$ 32.6/27.9/39.5 30.4/29.0/40.6 -11.9
HHH Reject DPO+$\mathcal{P}_{retr}$ PT+$\mathcal{P}_{retr}$ 48.0/21.8/30.1 39.3/25.8/34.9 +14.4
DPO+$\mathcal{P}_{gold}$ PT+$\mathcal{P}_{gold}$ 50.8/20.7/28.5 43.0/23.4/33.6 +20.2
Table 15: Comparison of personalization abilities of DPO and $\textsc{PT}_{\textsc{dpo}}$ when using the full Anthropic HHH and BT datasets. $\textsc{PT}_{\textsc{dpo}}$ still improves personalization on the rejected personas, but is slightly worse on the chosen personas. This is likely because DPO trained on chosen responses can already generate responses that tailor to chosen personas.
BeaverTails Anthropic HHH Mnemonic
Table 16: Win, tie, and loss rates of generation methods (FS, SFT, DPO) with and without personas $\mathcal{P}$ in pairwise comparisons from the Prometheus judge when using the filtered Anthropic HHH dataset. Our results are still strong compared to Table 2.